Test of cmmi10 on August 27, 2023 at 1737

´ 0  ´ 1  ´ 2  ´ 3  ´ 4  ´ 5  ´ 6  ´ 7
´ 00x  Γ  Δ  Θ  Λ  Ξ  Π  Σ  Υ ˝ 0x ´ 01x  Φ  Ψ  Ω  α  β  γ  δ  ϵ
´ 02x  ζ  η  θ  ι  κ  λ  μ  ν ˝ 1x ´ 03x  ξ  π  ρ  σ  τ  υ  ϕ  χ
´ 04x  ψ  ω  ε  ϑ  ϖ  ϱ  ς  φ ˝ 2x ´ 05x  ↼  ↽  ⇀  ⇁  ▷  ◁
´ 06x  0  1  2  3  4  5  6  7 ˝ 3x ´ 07x  8  9  .  ,  <  /  >  ⋆
´ 10x  ∂  A  B  C  D  E  F  G ˝ 4x ´ 11x  H  I  J  K  L  M  N  O
´ 12x  P  Q  R  S  T  U  V  W ˝ 5x ´ 13x  X  Y  Z  ♭  ♮  ♯  ⌣  ⌢
´ 14x  ℓ  a  b  c  d  e  f  g ˝ 6x ´ 15x  h  i  j  k  l  m  n  o
´ 16x  p  q  r  s  t  u  v  w ˝ 7x ´ 17x  x  y  z  ı  ȷ  ℘

˝ 8       ˝ 9       ˝ A       ˝ B       ˝ C       ˝ D       ˝ E       ˝ F